OpenStudy (anonymous):

y varies inversely with x, and x = -10 when y = 60. Find the inverse variation equation and use it to find the value of y when x = 15. -75 -90 -40 -30

OpenStudy (anonymous):

inverse relationship is: y = k/x they are asking you to find the value of k, then use the value of k to find y when x = 15 (that sounds confusing) so to start, find k by plugging in the given values at the beginning of the problem: 60 = k/-10 -6 = k now you have y = -6/x you should be able to solve for y knowing x = 15 here

OpenStudy (anonymous):

actually, i did that wrong it should be y = k/x 60 = k/-10 -600 = k which gives you a final equation of y = -600/x knowing x = 15 here, you should be able to find the correct answer